Use the inequality 5cos(2x)≤0

5cos⁡⁡(2x)≤0 on [−π,π]

[−π,π] to answer the following questions.

Part A: What are the zeros to the nearest hundredth?

Part B: What is the solution set for the inequality?

Select all correct answers for Part A, and select one answer for Part B. Use π≅3.14

π≅3.14 if needed.


A: x= −1/2π

A: x= −3/4π

A: x= 0π

A: x= −1/4π

A: x= 1/4π

A: x= 3/4π

A: x=−1/2π


B: (−3/4π,−1/4π)∪(1/4π,3/4π)

B: [−1/2π,0π]∪[0π,1/2π]

B: [−1/2π,1/2π]

B: [−3/4π,−1/4π]∪[1/4π,3/4π]

B: (−∞,−3/4π]∪[−1/2π,1/2π]∪[3/4π,∞)


1
Expert's answer
2021-06-12T04:48:30-0400
"5\cos(2x)\leq 0=>\cos(2x)\leq 0"

If "x\in[-\pi, \pi]," then "2x\in[-2\pi, 2\pi]"


"2x\in[-\dfrac{3\pi}{2}, -\dfrac{\pi}{2}]\cup[\dfrac{\pi}{2}, \dfrac{3\pi}{2}]"

"x\in[-\dfrac{3\pi}{4}, -\dfrac{\pi}{4}]\cup[\dfrac{\pi}{4}, \dfrac{3\pi}{4}]"

"5\cos(2x)= 0"

"=>x_1=-\dfrac{3\pi}{4}, x_2=-\dfrac{\pi}{4}, x_3=\dfrac{\pi}{4}, x_4=\dfrac{3\pi}{4}"


"=>x_1\approx-2.36, x_2=-0.79, x_3=0.79, x_4=2.36"


A: x= −3/4π

"\approx-2.36"


A: x= −1/4π

"\approx-0.79"


A: x= 1/4π

"\approx-0.79"


A: x= 3/4π

"\approx2.36"


B: [−3/4π,−1/4π]∪[1/4π,3/4π]